OpenStudy (anonymous):

Divide £4320 in the ratio 6:3:9? (PLZZZ HELLPPP ILL DOO GOOD!!)

OpenStudy (amistre64):

\[\frac{4320}{6:3:9}\]

OpenStudy (amistre64):

in other words: \[\frac{4320}{6}:\frac{4320}{3}:\frac{4320}{9}\]

OpenStudy (anonymous):

????? sooozzz but i am confuseedd plz explain moreee

OpenStudy (amistre64):

you take the given number, and divide it out

OpenStudy (anonymous):

okaiii so for the numberrrr divide it by each ratio (6:3:9) which will give u ur answeerrr?

OpenStudy (amistre64):

i believe so, but let me dbl chk that real quick

OpenStudy (anonymous):

thnxxxx x

OpenStudy (amistre64):

my thinking is off a bit; when we add the divisions, they should still equal 4320 and they aint

OpenStudy (anonymous):

ohhh?? well i am really stuck on thiss one it is quite hard for a person like me!!?

OpenStudy (amistre64):

i think then if we add the rations that will give us the total amount to divide by; then multiply that by the ratios

OpenStudy (amistre64):

that one works :)

OpenStudy (amistre64):

\[\frac{4320}{6+3+9}=1\ ratio\] \[\frac{4320(6)}{18}=6\ ratios\] \[\frac{4320(3)}{18}=3\ ratios\] \[\frac{4320(9)}{18}=9\ ratios\]

OpenStudy (anonymous):

okaiii thnxxx

OpenStudy (amistre64):

\[\frac{4320}{6+3+9}=240=1\ ratio\]

OpenStudy (anonymous):

youu r clevvverrrrr loolz

OpenStudy (amistre64):

does that make better sense?

OpenStudy (amistre64):

:) lol

OpenStudy (anonymous):

yuppp so the answerr issss 240??

OpenStudy (amistre64):

almost. 1 ratio is equal to 240; we need to count them up to get it into the ratios of 6:3:9 soo 240(6) : 240(3) : 240(9) and multiply those out should be right

OpenStudy (anonymous):

okaiiiiii soo for my question i should write down......

OpenStudy (anonymous):

sozzzz i am justt really stupid at math in the 2nd top group :s

OpenStudy (amistre64):

i trust you can multiply tho

OpenStudy (anonymous):

multiply 240 by 3?

OpenStudy (amistre64):

yes

OpenStudy (anonymous):

okaiii so that isss\[240\times3=720\]

OpenStudy (amistre64):

good, that gives us: 240(6) : 720 : 240(9) now multiply: 240 by 6 and again; 240 by 9

OpenStudy (anonymous):

1440 (x by 6) 1260 (x by 9)

OpenStudy (amistre64):

your "by 9" is either off, or typoed

OpenStudy (amistre64):

1440 : 720 : 2160 ill assume typoed :)

OpenStudy (anonymous):

soozzzzzz typed it in wroonggg

OpenStudy (amistre64):

thats your answer them

OpenStudy (anonymous):

sooo the answer isss : 1440 : 720 : 2160

OpenStudy (amistre64):

thats correct :)

OpenStudy (anonymous):

thnxx a loottt toook a long tymee but thnxxx !! ill add u as a fan"!

OpenStudy (amistre64):

youre welcome ;) and good luck with it all

OpenStudy (anonymous):

thnxxxx